OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12.

OpenStudy (bahrom7893):

subtract 4x from both sides. What do u get?

OpenStudy (bahrom7893):

4x + 3y - 4x = -12-4x Simplify that expression.

OpenStudy (anonymous):

3y=-12-4x y=-4-4/3x y=-4/3x-4

OpenStudy (bahrom7893):

darn it shiroh, don't just post the solns.. how are they gonna learn?

OpenStudy (anonymous):

But i did it step by step... D:<

OpenStudy (bahrom7893):

lol u think ppl are gonna read the steps. lol only one in a 100 will. Gage, it doesn't matter if ur school is virtual and u're doing it over spring break. U should still learn.

OpenStudy (anonymous):

Im not on break yet...

OpenStudy (bahrom7893):

u told me u were, but whatever.

OpenStudy (swag):

yes

OpenStudy (anonymous):

No sorry, that wasn't me. And that wasn't the rite answer. Here are the choices A. y = -12 - 4/3 x B. y = -4 - 4/3 x C. y = -12 + 4x D. y = -4 + 4x

OpenStudy (anonymous):

If you give me the answer and explain it, i will be able to understand it.

OpenStudy (anonymous):

The answer is there... B

OpenStudy (anonymous):

I just switched it around and put the variable in front. It's easier to read if the variables are in the front